(c) 

The forecast of a product for the first week of March was 200 units, whereas the actual 

demand turned out to be 220 units. 

i. 

Find the forecast for the second week of March by assuming the smoothing constant 

( ) α

as 0.35.      

ii. 

Find the forecast for the third week of March if the actual demand of the second 

week is 210 units.     

       


1
Expert's answer
2021-09-12T23:52:38-0400

"\text{}" "\text{Let D = Demand and F = previous forecast. The forecast for the second week is given }\\
\text{by}\\
\alpha(D-F)+F\\
\text{Hence, the forecast is}\\
0.35(220-200)+200=207\\
2. \text{ The forecast for the third week is;}\\
0.35(210-207) + 207 = 208.5"